OpenStudy (anonymous):

The series 2 + 4 + 8 + 16 + 32 + . . . is: arithmetic geometric both neither

OpenStudy (zzr0ck3r):

4/2 = 8/4 = 16/8 = 32/16 it is geometric

OpenStudy (secret66):

geometric

OpenStudy (shadowlegendx):

I was gonna say Bacon...but I guess that is wrong ;-;

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

And The series 6 + 6 + 6+ 6 + 6 is arithmetic? right?

OpenStudy (zzr0ck3r):

yes, the difference is constant as n runs

OpenStudy (anonymous):

it is arithmetic when last term is given

OpenStudy (anonymous):

so its neither i think becoz if it has last term we say both .if not neither